How many ounces of a 25% alcohol solution must be mixed with 10 ounces of a 30% solution to make a 26% alcohol solution
Answer by lwsshak3(11628)

How many ounces of a 25% alcohol solution must be mixed with 10 ounces of a 30% solution to make a 26% alcohol solution
***
let x=amt of 25% solution must be mixed
x+10=amt of mixed 26% solution
..
25%x+30%*10=26%(x+10)
.25x+3=.26x+2.6
.01x=.40
x=40
How many ounces of a 25% alcohol solution must be mixed? 40